OpenStudy (anonymous):

How do you use the Substitution Method to solve y=x^2 and y=x+6 in terms of finding the larger and smaller x value?

OpenStudy (anonymous):

y=x^2 and y=x+6 hence (x+6)=x^2 or x^2 -x-6=0 or x^2-3x+2x-6=0 or x(x-3)+2(x-3)=0 or (x-3)(x+2)=0 hence x=3 or x=-2 when x=3 y=3+6=9 and when x=-2 y=-2+6=4